OpenStudy (anonymous):

Ok this is a long one so please bear with me. I have this equation for y0 (it's the equilibrium distance for some sistem of springs and a mass). \[ 0 = m g +2k(1-\frac{l}{\sqrt{y_0^2+l^2}})y_0 \] So m is mass, k is spring constant and l is lenght of one spring. All well and good. I solved the equation for y_0 with Mathematica and got 4 solutions {y01,y02,y03,y04}. I am interested in the limiting value of y0 for m,k,g and l. I used Mathematica again to get: \[ m \rightarrow 0 \quad \left\{0,0,0,0\right\} \]

OpenStudy (anonymous):

(lol i ran out of space) \[ m \rightarrow \infty \quad \left\{ComplexInfinity, ComplexInfinity, -\infty, -\infty \right\} \] \[ g \rightarrow 0 \quad \left\{0,0,0,0\right\} \] \[ g \rightarrow \infty \quad \left\{ComplexInfinity, ComplexInfinity, -\infty, -\infty \right\} \] \[ k \rightarrow 0 \quad \left\{ComplexInfinity, ComplexInfinity, -\infty, -\infty \right\} \] \[ k \rightarrow \infty \quad \left\{0,0,0,0\right\} \] \[ l \rightarrow 0 \quad \left\{ DirectedInf\left[\frac{1}{2}(1-i \sqrt{3})\right],DirectedInf\left[\frac{1}{2}1+i \sqrt{3}\right],-\infty ,ComplexInfinity \right\} \] \[ l \rightarrow \infty \quad \left\{0,0,-\frac{g m}{2 k},-\frac{g m}{2 k}\right\} \] All the limits except the second last one make sense to me. Bigger mass or stronger g makes the ball drop lower. And stronger spring (k) holds the ball higher up (closer to 0). (By looking at the limits I deduced that solution number 3 is the right one). Finally my question is why when I nondimensionalize the equation with: \[ \gamma = \frac{y}{y_c} \] \[ y_c = \frac{mg}{2k} \] \[ \tau = \frac{t}{t_c} \] \[ t_c = \sqrt{\frac{m}{2k}} \] I get the equation, \[ (1-\frac{1}{\sqrt{(\frac{mg}{2kl})^2 \gamma_0^2 + 1}})\gamma_0 + 1 = 0 \] and define a new parameter, \[ \alpha = (\frac{mg}{2kl})^2 \] and look at the limiting values of alpha I get different results for gamma0??? I would think that a large alpha would mean stronger gravitational force therefore a lower equilibrium distance gamma0. And with a smaller alpha I would assume to get gamma0 = 0. But \[ \alpha \rightarrow 0 \quad \left\{ DirectedInf\left[\frac{1}{2}(1-i \sqrt{3})\right],DirectedInf\left[\frac{1}{2}1+i \sqrt{3}\right],-\infty ,-\frac{1}{2} \right\} \] \[ \alpha \rightarrow \infty \quad \left\{ 0,0,-1,-1 \right\} \] Meaning that a big alpha means y0 goes to -(mg)/(2k) and a small alpha means y0 goes to -inf. This is exactly the opposite of what I got before nondimensionalization... Someone please help me.

mathslover (mathslover):

@hartnn and @UnkleRhaukus

OpenStudy (anonymous):

@Callisto @UnkleRhaukus @ganeshie8 @AccessDenied @ghazi

OpenStudy (ghazi):

@shfreeman i appreciate your hard work :) ...

OpenStudy (anonymous):

Thank you. Can you help out a fellow hard worker? I would appreciate if someone could try the things I did and report on their results. So we can at least see that there is a mistake somewhere..

OpenStudy (ghazi):

i have to go through this...because i can see there is a mistake whilst you were doin nondimensionalization ...but i am not sure...so i'll take time thank you